OpenStudy (anonymous):

What is the y-intercept of the function that passes through the points (–3, 12) and (4, –2)?

OpenStudy (anonymous):

your first step would be to find the slope. Once you find the slope you can use the equation: (y-y1)=m(x-x1) where you plug in your point (either (-3,12) or (4, -2) ) into the (x1,y1) and then simplify into y=mx+b